OpenStudy (anonymous):

4 adults and 3 kids are seated in a row. In how many different ways can they be seated if an adult must occupy the first position in the row?

OpenStudy (anonymous):

@satellite73 can u help me out, and see if my answerrs are right?

OpenStudy (anonymous):

what abt this question? Private automobile license plates contain three letters followed by three numbers. If there are no other constraints, the number of plates that could be manufactured is

OpenStudy (anonymous):

is the answer 17576000

OpenStudy (anonymous):

@callisto can u help me out?

OpenStudy (callisto):

I'm not sure if it is correct... No. of ways = 4x 6x5x4x3x2x1 = ... The first one is 4 as one of the 4 adults must occupy the first position. So, 4 people can be chosen..

OpenStudy (callisto):

Sorry, I mean for the first question :|

OpenStudy (callisto):

For the second question, can the numbers/letters be used again?

OpenStudy (callisto):

(My English fails me :( )

OpenStudy (callisto):

If it can, then I think it is like this: No. of ways = 26^3 x 10^3 =.... 26 => 26 English letters 10 => 10 numbers (0, 1, 2, 3, ..., 7, 8, 9)

OpenStudy (anonymous):

@lgbasallote can you double check the first question?

OpenStudy (lgbasallote):

uhmm sorry..im not good with permutations and combinations :/

OpenStudy (anonymous):

@robtobey can you help me

OpenStudy (anonymous):

i think it is \(4\times 6!\) 4 choices for the first seat and then \(6!\) for the remaining 6

OpenStudy (anonymous):

oh what callisto said above

OpenStudy (anonymous):

second one is what callisto said as well both answers are right